OpenStudy (anonymous):

Find all solutions to the equation. 7 sin2x - 14 sin x + 2 = -5

OpenStudy (alekos):

let sinx = t and we get 7t^2 - 14t + 2 = -5 so what do you think comes next?

OpenStudy (anonymous):

I'm not sure, subtract 2?

OpenStudy (alekos):

No. add 5 to both sides, and we get?

OpenStudy (anonymous):

7t^2 - 14t + 7 ?

OpenStudy (alekos):

That's correct but it's equal to zero. So we have a quadratic equation! what do you think the next step might be?

OpenStudy (anonymous):

I can't seem to remember, can you help?

OpenStudy (alekos):

we have to simplify the equation and then solve for t 7t^2 - 14t + 7 = 0 can you see a common factor?

OpenStudy (anonymous):

divide by 7t?

OpenStudy (alekos):

that's close. Try dividing by 7. Now what does that become?

OpenStudy (anonymous):

t^2-2 = 0?

OpenStudy (alekos):

nice try, but not quite. Dividing everything by 7, We get t^2 - 2t + 1 = 0 Do you follow?

OpenStudy (shadowlegendx):

o_O

OpenStudy (anonymous):

Yes!

OpenStudy (alekos):

OK, now we have to factorise the quadratic. Any ideas?

OpenStudy (anonymous):

I think I figured it out, t = 1 right?

OpenStudy (alekos):

Well done!! So now we have t = 1 which means that sinx = 1 now we have to work out what values x can have. Any ideas?

OpenStudy (anonymous):

1? Haha

OpenStudy (alekos):

Very funny :) x = arcsin1 any ideas what that might be?

OpenStudy (anonymous):

No sorry, this isn't really my greatest subject

OpenStudy (alekos):

Oh, I see. Have you tried using a calculator or google?

OpenStudy (alekos):

We're looking for an angle that has a sine of 1

OpenStudy (anonymous):

Can you give me a hint?

OpenStudy (alekos):

OK, what is the sine of π/2 ?

OpenStudy (anonymous):

1?

OpenStudy (alekos):

Yes, that's right. So that means that x = arcsin 1 = π/2 But that's not all! Do you know why?

OpenStudy (anonymous):

No, so is 1 the answer?

OpenStudy (alekos):

No, the answer is x = π/2 + 2kπ where k = all integers

OpenStudy (anonymous):

thank you!

OpenStudy (alekos):

Because the answer repeats every 2π

OpenStudy (alekos):

No problem. Glad I could help